how do you solve 4m/3-m/6=7/2
Answer by tracyc0912(40)

4m/3-m/6=7/2 (multiply the whole thing by 6 to remove fractions)
8m-1m=21, 7m=21 (divide both sides by 7
m=3
Thanks!